OpenStudy (anonymous):

A chef is using a mixture of two brands of Italian dressing. The first brand contains 9% vinegar and the second brand contains 14% vinegar. The chef wants to make 400 milliliters of a dressing that is 11% vinegar. How much of each brand should she use?

OpenStudy (anonymous):

You can put this solution on YOUR website! I am a visual learner, so I try to draw a picture to help me understand a word problem. In a mixture problem, you are working with an amount(ml) and a concentration(%). I draw 3 lines ___+___=___ (I usually use boxes instead of lines to combine the #'s and variables) Let your first brand = x Let your second brand = y You do not need a variable for the product because you have an amount and a concentration already. Use the variables (x and y) to replace your unknowns. The unknowns are the amounts of the first and second brands. x+y=320 (this is your equation for the amounts) 0.08x +0.13y=.11(320) (this is your equation for the concentrations) We can use the substitution or elimination method to solve this. I will use the substitution method. I don't like using decimals, so I am going to move the decimal place two units to the right to make them whole numbers. This gives me: 8x+13y=11(320) Using substitution on the first equation, I put X on one side of the equation. This gives me x=(-y+320). I can now substitute (-y+320) for x in the second equation. 8(-y+320)+13y=3520 -8y+2560+13y=3520 Subtract 2560 from both sides and combine 13y and -8y. This gives you 5y=960. Divide both sides of the equation by 5. This gives you y=192. y=192 x=128 I got x by plugging y into equation #1. x+(192)=320. After subtracting 192 from 320, I get x= 128. To check your answer, always substitute x and y in for both equations. ANSWER: The chef should use 128 milliliters of the 8% vinegar brand and 192 milliliters of the 13% vinegar brand.

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

Can I have a medal or a fan or both please

OpenStudy (anonymous):

wait, how did get 8 and 13. The amounts are 9 and 14? I'm confused

OpenStudy (anonymous):

Um give me a minute to think

OpenStudy (anonymous):

one step ahead of you @Jannah☺️ I already fanned you both

OpenStudy (anonymous):

The unknowns are the amounts of the first and second brands. x+y=320 (this is your equation for the amounts) 0.08x +0.13y=.11(320) (this is your equation for the concentrations) We can use the substitution or elimination method to solve this.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

no, im still trying to figure out how you got 8 and 13 from 9 and 14....it was a pretty good example though

ganeshie8 (ganeshie8):

still stuck on this problem ?

OpenStudy (anonymous):

yes, still trying to figure out how 9 and 14 became 8 and 13

OpenStudy (anonymous):

wait I have it. they gave me an example. if I substitute using my formulas is would be: brand 1:240 ml brand 2: 160 ml is that right @ganeshie8 @ikram002p

OpenStudy (anonymous):

not formulas...I meant values.

ganeshie8 (ganeshie8):

this is easy problem. lets work it from SCRATCH, real quick ok ?

ganeshie8 (ganeshie8):

u wid me ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

say, he uses \(A\) mL from first brand, and \(B\) mL from second brand, then : \(A + B = 400 \) ---------------(1)

ganeshie8 (ganeshie8):

since he wants it to be of \(11\%\) vinegar : \(\large (9\%~ of ~A) + (14\%~ of ~B) = (11\%~ of ~400) \) ----------(2)

ganeshie8 (ganeshie8):

two equations, two unknowns. u can solve them.

ganeshie8 (ganeshie8):

let me knw if something doesnt make sense :)

OpenStudy (anonymous):

im with you

OpenStudy (anonymous):

I still have brand 1: 240 brand 2: 160 @ganeshie8

ganeshie8 (ganeshie8):

let me check

ganeshie8 (ganeshie8):

equation (2) becomes : \(\large 0.09A + 0.14B = 0.11*400\)

ganeshie8 (ganeshie8):

solving them both gives : \(\large A = 240,~~ B=160\) http://www.wolframalpha.com/input/?i=A%2BB%3D400%2C+.09A+%2B+.14B%3D.11*400

ganeshie8 (ganeshie8):

Good job !! @AwesomeAries

OpenStudy (anonymous):

Thank you @ganeshie8

ganeshie8 (ganeshie8):

np :)